OpenStudy (anonymous):

find the ordered pair for the ertex. y+x^2-14x+51

OpenStudy (anonymous):

y=x^2-14x+51

OpenStudy (shadowfiend):

Given \(ax^2 + bx + c\) ^^

OpenStudy (anonymous):

sorry gents, just want the answer, this is my last question for this entire class

OpenStudy (shadowfiend):

Understandable, but not the way I roll, sorry ;) Someone else'll have to sweep in.

OpenStudy (anonymous):

no problem shadow, take care

OpenStudy (sasogeek):

you gotta do it urself :)

OpenStudy (anonymous):

bye saso, there is a nice lady below me who needs help, move along

OpenStudy (sasogeek):

seriously u don't tell me where to look or who to help :) thank you for ur concern :D

OpenStudy (anonymous):

rude

OpenStudy (shadowfiend):

As a side note, you may want to try Wolfram Alpha.

OpenStudy (anonymous):

i have tried it, too complicated

OpenStudy (anonymous):

(7, 2)

OpenStudy (shadowfiend):

Really? I just typed “vertex of y=x^2-14x+51” into it and it gave me the answer ;)

OpenStudy (anonymous):

ah, 0,-7 thank you iran

OpenStudy (sasogeek):

if you want answeres... wolfram alpha is there for ya... if u come here, get ready to do part of the work, i'm not ur calculator ;)

OpenStudy (anonymous):

you can also use a TI-89, solve for the derivative equated to 0, and then evaluate the original expression at x = the solution for the derivative equal to 0

OpenStudy (anonymous):

no one asked you to be saso, in fact, i don't remember saying anything about you in my questions

OpenStudy (shadowfiend):

sasogeek—I think you've made your point. bhruifafn—likewise, I think you've made your apathy evident.

myininaya (myininaya):

Ok I will show you how to derive a simple formula for finding the vertex \[ax^2+bx+c=a(x^2+\frac{b}{a}x)+c=a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)+c-a(\frac{b}{2a})^2\] \[=a(x+\frac{b}{2a})^2+c-a \frac{b^2}{4a^2}\] \[=a(x+\frac{b}{2a})^2+c-\frac{b^2}{4a}=a(x+\frac{b}{2a})^2+\frac{4ac-b^2}{4a}\] Vertex is \[(-\frac{b}{2a},\frac{4ac-b^2}{4a})\]

OpenStudy (anonymous):

lol thank you mother shadow

OpenStudy (shadowfiend):

Mm. No problem. Just remember this mother has suspension powers. Though it sounds like you won't be needing help anymore after this fella.

OpenStudy (anonymous):

mother shadow lol

OpenStudy (anonymous):

suspend away mother dearest

OpenStudy (anonymous):

this is it for me :) thanks to everyone who helped!!

OpenStudy (anonymous):

7,2 or 0,7?

OpenStudy (shadowfiend):

(7, 2)

OpenStudy (anonymous):

thank you shadowfiend. the help was appreciated, if you have so much power you should suggest it be known that the person asking the question will be required to do work, not just ask for answers

OpenStudy (anonymous):

and that harassment is rampant

OpenStudy (shadowfiend):

Harassment is something we're working on. We're also working on reining in some of our more... Erm... Emphatic helpers, and making it clear that if they don't want to give out answers they should move on rather than harass.

myininaya (myininaya):

\[(-\frac{-14}{2(1)},\frac{4(1)(51)-(-14)^2}{4(1)})=(7,2)\] you could get what shadow got just by plugin the answer into this formula for the vertex i gave you

OpenStudy (anonymous):

thank you myininaya

OpenStudy (anonymous):

and shadow that sounds wonderful. take care. goodbye openstudy, hopefully forever, but possibly not. time will tell :). mwuah. it's been real.

OpenStudy (anonymous):

come on dont leave we'll all miss you

OpenStudy (anonymous):

hmm... what's the Aesop we can draw from all this?

OpenStudy (shadowfiend):

The moral of the story is, just because you see a hot girl avatar... Doesn't necessarily mean you're dealing with a girl ;)

OpenStudy (anonymous):

bingo

myininaya (myininaya):

lol what?

OpenStudy (anonymous):

Well, ill miss the banter, is what i meant

OpenStudy (anonymous):

study complete. you get much more help if you put up a pic of a hot girl, and sudo act like one

OpenStudy (shadowfiend):

Hehehe. You would get along swell with our CEO.

OpenStudy (anonymous):

you have all helped my senior thesis. thank you for participating in my abnormal psych study.

myininaya (myininaya):

I don't go where the hot girls are.

OpenStudy (anonymous):

I should have left my girl avatar on :(

myininaya (myininaya):

I just go to a question I want to answer.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

you didn't answer my pseudo-inverse question :(

OpenStudy (anonymous):

not what i have found out, you are 82% more likely to answer my question because of the arousing pic, than the normal owl

myininaya (myininaya):

I looked at it, but that requires more brain power than I have right now.

OpenStudy (anonymous):

and it took someone in the know to out me.

OpenStudy (anonymous):

let me see your data on that

OpenStudy (anonymous):

hmm maybe JamesJ could help me with it when he has time.

OpenStudy (anonymous):

it's basic empathy

OpenStudy (anonymous):

no one helps the dirty bum, but the damsel in distress?

OpenStudy (anonymous):

and i wondered why no one was helping me

OpenStudy (shadowfiend):

Indeed. I must say, too, it was a well-picked username. Seemed like a perfectly viable username of Arabic origin, with a perfectly plausible Arabic avatar.

OpenStudy (anonymous):

ching ching, thank you shadow

OpenStudy (anonymous):

but i couldn't beat you

OpenStudy (anonymous):

wanna know the best part shadow?

OpenStudy (shadowfiend):

I have insider info. Doesn't count ;)

myininaya (myininaya):

I don't even know what to type. I'm so confused.

myininaya (myininaya):

I was following the crowd not the pretty picture.

OpenStudy (anonymous):

Try me next shadow? what do you know about me?

OpenStudy (anonymous):

if you take bhruifafn and take every other letter, you get b r i a n h u f f my real name, male origin.

OpenStudy (anonymous):

Shadow knows all, tbh.

OpenStudy (anonymous):

i am

OpenStudy (anonymous):

it worked, the response time to my question was cut almost in half

OpenStudy (shadowfiend):

The psychologista 8-) A'ight, time to go fix some bugs hehe. Thanks for the entertainment.

OpenStudy (anonymous):

thanks shadow. take care man.

OpenStudy (chris):

It was never about the golden goose

OpenStudy (chris):

I presume

OpenStudy (anonymous):

no i didn't, my class/thesis is complete

OpenStudy (chris):

Brian - this is awesome stuff. Would you mind sending me a copy of the thesis once you've turned it in and gotten your A :)

OpenStudy (anonymous):

if i could give any advice, it would be make a female name and female avatar, people jump all over your questions

OpenStudy (anonymous):

i suppose i cold do that chris

OpenStudy (chris):

Well, not so awesome in that you tricked your fellow peers - but this speaks of greater biases, even within an altruistic community

myininaya (myininaya):

Don't include me in your statistics. I was just following the crowd. serious stuff. How do you know the others weren't doing the same?

OpenStudy (anonymous):

well i agree, but the people in the know (shadowfiend) saw i was a male, because i put my sex as male in beginning

OpenStudy (anonymous):

i don't know that myininaya. it's also about the level of help given, when i was male i was expected to know more, when female people walked me through more and more willing to just give me the answer.

OpenStudy (anonymous):

Here's an idea for your next thesis:subjects with names that are egregiously long and complicated, and with black male avatars do not get their questions answered as adequately as the female ones do :-P

OpenStudy (anonymous):

nope, just fun om

OpenStudy (anonymous):

lol i love it ag

OpenStudy (shadowfiend):

As a side note, I didn't use any male/female data. You said something (I forget what at this point) that made me rethink, and I have access to your email. That was sufficient in this case.

myininaya (myininaya):

I have given tons of people detailed answers without looking at their pic.

OpenStudy (anonymous):

the ends justify the means shadow

OpenStudy (anonymous):

well myininaya you are the outlier, always one or two statistical misfits

OpenStudy (chris):

myin - everyone knows you help everybody without bias =). don't worry about anyone drawing conclusions about individuals - you weigh equally on both sides.

myininaya (myininaya):

thank you thats all i wanted to hear :)

OpenStudy (anonymous):

lol i am not including any names in this at all, just numbers

OpenStudy (anonymous):

can i see your data? or is it priviate?

OpenStudy (anonymous):

even if i was, it's like taking a photo in public, this is a public domain and anything you say/do here can be recorded

OpenStudy (anonymous):

private for now lagrange

OpenStudy (anonymous):

here meaning "internet"

OpenStudy (chris):

Brian - please do not share any personally identifiable information/analysis - generally studies do not allow so (at least for journals). Aggregate analysis that you've mentioned is fair game.

OpenStudy (chris):

which I understand was your intent

OpenStudy (anonymous):

exactly chris

OpenStudy (shadowfiend):

The admins do, yes.

OpenStudy (anonymous):

i have some morals

OpenStudy (anonymous):

up to the point to get me an a